OpenStudy (anonymous):

Write a quadratic variation equation if g(x) varies directly with x^2, and g(x) = 98 when x = 7. A. g(x)= 7x^2 B. g(x)= 1/2x^2 C. g(x)= 14x^2 D. g(x)= 2x^2

OpenStudy (anonymous):

you asked almost the same question before, just there different numbers

OpenStudy (anonymous):

g(x)=k*x^2 , substitute g(x)=98 to find "k". After use that equation and substitute for x=7

OpenStudy (anonymous):

98=k*7^2

OpenStudy (anonymous):

answer is D